OpenStudy (anonymous):

Please help Solve 3(tan^4(x))+4(tan^2(x))-1=0 I know there are multiple situations but I was wondering how to solve this algebraically like a polynomial without graphing

rebeccaxhawaii (rebeccaxhawaii):

http://www.calculators.org/math/algebra.php

OpenStudy (anonymous):

Not situations, solutions

rebeccaxhawaii (rebeccaxhawaii):

plug it in here for great explanation be sure to plug in

OpenStudy (anonymous):

Umm actually that doesn't work because it asks for an upgrade for step by step solutions

OpenStudy (tkhunny):

There should be no plugging in. Solve this: \(3x^{2}+4x-1 = 0\)

OpenStudy (anonymous):

\[x=\frac{ -2\pm \sqrt{7} }{ 3 }\]

OpenStudy (anonymous):

That's what I got Would that squared be equivalent to using the 4th power instead of the second like we did?

OpenStudy (tkhunny):

Okay, now write \(\tan^{2}(x) = \dfrac{-2\pm\sqrt{7}}{3}\), think about why I did that, and you're almost done.

OpenStudy (anonymous):

Ohhhhh Wow that makes sense, thanks a lot!

OpenStudy (anonymous):

You would take the square root of that expression and get \[\tan x=\sqrt{\frac{ -2\pm \sqrt{7} }{ 3 }}\] right?

OpenStudy (tkhunny):

No quite. Add another \(\pm\)

OpenStudy (anonymous):

Oh right in front of the square root Thank you!